Rylee88:

Simplify this: (10*3)^19 I don't understand 😭

Rylee88:

*^19 is an exponent*

spicycautionramen:

well when they say simplyfy we have to get ride of the exponet right? or no

Rylee88:

Yes, but how so?

spicycautionramen:

so we use pemdas?

Rylee88:

OH, yes!

Rylee88:

So 10*3 right? which would be 30^19

spicycautionramen:

yes

Rylee88:

Then 30, 19 TIMES??

spicycautionramen:

yeah which will give you...?

Rylee88:

Sorry lol I had to switch my class, I have to do this? (30*30*30*30*30*30*30*30*30*30*30*30*30*30*30*30*30*30*30)

spicycautionramen:

uh this is what i got when i did it does it look about right?

1 attachment
Rylee88:

Yeah, that's what I got too.

Rylee88:

But that's definitely wrong for this question.

UhhhKhakis:

Is there a difference between * and ^?

spicycautionramen:

@uhhhkhakis wrote:
Is there a difference between * and ^?
^ is for an exponet and *is multiply

spicycautionramen:

and is it suposed to be * and not /?

UhhhKhakis:

@spicycautionramen wrote:
@uhhhkhakis wrote:
Is there a difference between * and ^?
^ is for an exponet and *is multiply
So why would 30 to the power of 19 be an answer with numbers, symbols and letters.

UhhhKhakis:

Awfully weird.

Rylee88:

1 attachment
spicycautionramen:

@uhhhkhakis wrote:
@spicycautionramen wrote:
@uhhhkhakis wrote:
Is there a difference between * and ^?
^ is for an exponet and *is multiply
So why would 30 to the power of 19 be an answer with numbers, symbols and letters.
well if i rembe right when its to the power of something you have to do 30*19 30*19 until theres 19 of 30*19s

Rylee88:

Yes you're correct, but that can't be the answer.

UhhhKhakis:

So basically you have to find (or remember) a law of exponents to simplify the expression. Remember any? (and stop with the names)

Rylee88:

It literally says it on the ss, read.

Rylee88:

Power of Products

UhhhKhakis:

Oh okay then.

UhhhKhakis:

1 attachment
Rylee88:

I know, I'm the one who put Power of Products

Rylee88:

I have to SIMPLIFY (10*3)^19

Rylee88:

@spicycautionramen do you know how to do it? πŸ€”

spicycautionramen:

@rylee88 wrote:
@spicycautionramen do you know how to do it? πŸ€”
im thinking ot doing it the oposit way or backwards to see if i can get the same thing

Rylee88:

Hm, so how would be do that?

UhhhKhakis:

You have to follow the Order of Operations tho

spicycautionramen:

oposite would be sum like 10/3*19/1 those are fractions being multiplyed

Rylee88:

But that's not the question.

spicycautionramen:

i know hut im looking for other ways to do it cuz idk i walked in late on these notes

Rylee88:

Ahh.

Rylee88:

How would you even get that? Also don't give out direct answers.

spicycautionramen:

does 63 1/3 seem more reaasonable?

Rylee88:

No one knows?

Extrinix:

Using the power of products would make it something like this: \(\sf{10^{19}\times 3^{19}}\) Given that \(\sf{(ab)^{x}=a^x b^x}\)

Extrinix:

As for simplifying it more than that, if you’re also using exponent rules, the equation would be something like (\(\sf{(10 \times 3) ^{19+19}}\)

Rylee88:

I'll be right back ( I have to change classes)

Extrinix:

πŸ‘

Rylee88:

I'm back c:

Extrinix:

Welcome back

Rylee88:

Thank you.

Rylee88:

@extrinix wrote:
As for simplifying it more than that, if you’re also using exponent rules, the equation would be something like (\(\sf{(10 \times 3) ^{19+19}}\)

Rylee88:

19+19=38

Rylee88:

10*3=30

Rylee88:

THE ANSWER: 10^19 * 3^19 BRUH

Ferrari:

I love meth

Ferrari:

I mean math i said math math